OpenStudy (kewlgeek555):

I Just Need to Know How to Simplify This:

OpenStudy (kewlgeek555):

\[(3^{\frac{ 2 }{ 3 }})^{\frac{ 6 }{ 5 }}\]That is the expression I would like to simplify. I just want to know HOW to do it, I want the steps, not the answer.

OpenStudy (jdoe0001):

\(\huge (3^{\frac{ 2 }{ 3 }})^{\frac{ 6 }{ 5 }}\implies 3^{\frac{ 2 }{ 3 }\cdot \frac{ 6 }{ 5 }} \)

OpenStudy (kewlgeek555):

Okay. So that is the...I don't know what property of exponent that is...

OpenStudy (kewlgeek555):

Bur I can do that.

OpenStudy (kewlgeek555):

Soit could be...\[3^{\frac{ 4 }{ 5 }}\]or\[3^{0.8}\]since they are the same thing...

OpenStudy (kewlgeek555):

Thanks @jdoe0001

OpenStudy (jdoe0001):

yw